A.   Now, assuming the economy is open with government (G) participation and  external trade which is summarized as follows; export(X)= 100-0.10Y, import(M)=50, G=100, Taxes(T)= 100 and C, I, MS, Mt, and Mz the same as defined in (a) above. Calculate;

   i.           The equilibrium income and interest rate in this new economy.         

  ii.           The level of C, I, Mt, and Mz when the economy is in equilibrium      

 


1
Expert's answer
2021-07-30T14:26:01-0400

Solution


"IS= Y= C+I+G+T+(X-M)"


"Y=102+0.7Y+150+100+(100-0.10Y-50)"


"Y=1255-250r (IS Curve)"


"LM=437.5 +500r"


At equilibrium

IS=LM


"1255-250r=437.5 +500r"


solve for r

r= 1.09 is equilibrium rate of interest


"Y=1255-250r"


"Y= 982.5" is equilibrium income


2. "C=102+0.7Y"


"C= 102+0.7\times982.5" "=789.75"


"I=150-100r"


"I=150-100\times1.09=41"


"Mt=0.4Y"


"Mt= 0.4 \times 982.5" "=393"


"Mz=125-200r"


"Mz=125-200\times1.09=-93"